the sum of two numbers is 90. The larger number is twice he smaller number. What are the two numbers?
Answer by bucky(2100)

Let L represent the larger of the two numbers and let S represent the smaller of the two numbers.
.
The sum of these two numbers is 90. In equation form that is written as:
.
L + S = 90
.
Next the problem tells you that the Larger number is twice the smaller. This means that 2 times
the smaller number (or 2S) is equal to the larger number (or L). In equation form this is:
.
L = 2S
.
We can take the right side of this second equation and substitute it into the first equation
in place of L because 2S is equal to L. When we substitute 2S for L in the first equation, the
first equation becomes:
.
2S + S = 90
.
Add the terms on the left side and you have:
.
3S = 90
.
Solve for S by dividing both sides of the equation by 3, the multiplier of S, and you have:
.
S = 90/3 = 30
.
The smaller number is 30. And since the total of the two numbers is to be 90 we can find the
larger of the two numbers by subtracting 30 from 90. The result is that the larger, L, is
equal to 90 - 30 = 60.
.
Note that the two numbers are 60 and 30.
.
The sum of these two numbers is 60 + 30 = 90
.
The larger number (60) is twice the size of the smaller number (30) just as the problem said
that it was to be.
.
So our answer checks out.
.
Hope this helps you to understand the problem and how to work it.
.